# A Comprehensive Survey on Automatic Text Summarization with Exploration of LLM-Based Methods

### v2

@@ 6.2.2 Fine-tuning LLMs for ATS
Internal efficient-parameters fine-tuning: This approach involves freezing most of the LLM’s parameters and updating only a few layers, such as the embedding or output layers. [69] enhanced model efficiency by replacing the self-attention layers in the encoder with disentangled attention layers, representing each word with two vectors to encode its content and position. Similarly, [257] explored efficient fine-tuning strategies for large language models in radiology report summarization, utilizing domain-specific pre-training while updating only 0.32% of the model’s parameters.
@@
External adapters fine-tuning: Training external adapters allows models to learn the distribution of target data without modifying the original parameters of the LLM. [258] proposed an efficient few-shot method using adapters, where the adapters were pre-trained on a large corpus and fine-tuned on a smaller human-annotated dataset. Similarly, [259] introduced a contrastive learning approach for supervised ATS, aiming to maximize the similarity between a document, its gold-standard summary, and model-generated summaries. Additionally, [260] combined encoder and decoder foundation models into a single model, AdaMo, and implemented adaptive knowledge transfer techniques such as continuous pretraining and intermediate fine-tuning, along with task-specific designs for sequence-to-sequence learning.
6.2.3 Knowledge Distillation from LLMs for ATS
@@
A knowledge distillation system consists of three key components: knowledge, a distillation algorithm, and a teacher-student architecture [275]. In this process, a smaller student model is supervised by a larger teacher model [276]. For ATS, this typically involves an offline distillation process where an LLM serves as a fixed teacher model, with its outputs providing additional supervision to train a smaller student model. This approach enables the student model to inherit the LLM’s robust summarization capabilities, making it particularly useful in scenarios with limited computational resources or strict data privacy requirements. For example, [228] improved the quality of aspect-triple rationales and summaries through a dual scoring system, followed by a curriculum-learning strategy to train a smaller local model on ATS tasks. Similarly, [261] addressed low-resource cross-lingual summarization using mBART for incremental training and reinforcement learning to optimize discrete cues, while [243] proposed a three-step method for adapting smaller models to summarize forum discussions by sampling from a large corpus, retrieving annotated prompts, and filtering low-quality data.
@@ 7.1 Metrics based on Term Overlap
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) [279] is a widely used metric for ATS evaluation, measuring overlap between candidate summaries ($C$) and reference summaries ($S$).
ROUGE-N calculates n-gram overlap as:
@@
METEOR (Metric for Evaluation of Translation with Explicit ORdering) [281] was developed to overcome the rigidity of BLEU, which relies on exact n-gram matches. METEOR evaluates summaries by aligning words between a candidate and reference summary using flexible matching techniques, including synonyms, stemming, and paraphrases, facilitated by resources like WordNet. This allows METEOR to account for linguistic variations and semantic similarities that BLEU cannot capture. Additionally, METEOR differentiates between function words (e.g., ”the,” ”and”) and content words (e.g., ”run,” ”house”), giving more weight to content words to better reflect meaning. By incorporating these features, METEOR provides a more nuanced and flexible evaluation metric for summarization tasks.
@@ 7.2 Similarity-based ATS Metrics
Similarity-based metrics assess summary quality by measuring the semantic similarity between the candidate summaries with the reference summaries.
BERTScore and BARTScore [283, 285] evaluates text quality by computing a similarity score between each token in the candidate summary and each token in the reference summary, using contextual embeddings from pre-trained language models like BERT or BART. This approach captures semantic meaning and contextual relationships, offering a more flexible and effective evaluation compared to traditional n-gram-based metrics.
BLEURT [285] builds on BERT embeddings by incorporating additional fine-tuning on human-annotated data and synthetic examples. This fine-tuning process allows BLEURT to explicitly model text quality judgments, aligning more effectively with task-specific requirements such as fluency, coherence, and relevance. Comparing to BERTScore, BLEURT provides a more comprehensive evaluation by integrating various aspects of text quality.
@@
[…]
@@
FactCC [171] takes a source document and a summary statement as input and outputs the fact consistency label for that summary statement. FactCC is a weakly-supervised model proposed to verify factual consistency in summaries by predicting consistency, extracting supporting spans, and identifying inconsistent spans, which outperforms strongly supervised models and aids human verification. By modeling different types of fact consistency errors, the method achieves performance improvement in fact consistency evaluation.
@@ 7.3 LLM-based ATS Metrics
Traditional metrics for evaluating summary quality often focus on specific dimensions, such as lexical overlap or semantic similarity, but they fail to capture the full complexity of summarization tasks. Large Language Models (LLMs) have introduced a paradigm shift in evaluation by leveraging their deep contextual understanding and semantic reasoning capabilities, enabling more comprehensive assessments.
Writing Quality Metrics: Prompt-based LLM metrics are increasingly used to assess the overall writing quality of summaries across multiple dimensions. [292] explored ChatGPT’s ability to replicate human evaluation methods such as Likert scoring, pairwise comparison, Pyramid, and binary factuality, outperforming traditional metrics on some datasets. [293] used GPT-3 with in-context examples to evaluate dimensions like consistency, relevance, fluency, and coherence. [294] introduced a role-player prompting mechanism to generate judgments from multiple roles, achieving high consistency with human evaluations. [295] proposed BOOOOKSCORE, which used GPT-4 prompts to identify issues like causal omissions, salience, and duplication.
Faithfulness and Factuality Metrics: Faithfulness and factuality metrics evaluate whether summaries accurately represent the source text. [296] examined ChatGPT’s ability to detect factual inconsistencies in zero-shot settings across tasks like binary entailment inference, summary ranking, and consistency rating, demonstrating strong performance compared to prior metrics. [297] introduced FIB (Factual Inconsistency Benchmark), showing that LLMs typically assign higher scores to factually consistent summaries. [298] proposed TrueTeacher, which uses LLMs to annotate model-generated summaries for factual consistency and generate synthetic data. [299] developed FFLM, a zero-shot faithfulness evaluation method leveraging probability changes, outperforming or matching ChatGPT in detecting inconsistencies and assessing faithfulness.

### v3

@@ 7.3.1 Internal efficient-parameters fine-tuning
This approach involves freezing most of the LLM’s parameters and updating only a few layers, such as the embedding or output layers. [250] addressed the challenge of summarizing lengthy biomedical literature for lay audiences by proposing an extract-then-summarize framework leveraging LLMs. The researchers implemented a fine-tuned GPT-3.5 model within this framework to enhance the system’s ability to generate accessible summaries tailored to non-expert readers. [67] enhanced model efficiency by replacing the self-attention layers in the encoder with disentangled attention layers, representing each word with two vectors to encode its content and position. Similarly, [249] explored efficient fine-tuning strategies for large language models in radiology report summarization, utilizing domain-specific pre-training while updating only 0.32% of the model’s parameters.
Reinforcement Learning from Human Feedback (RLHF) [251] offers a powerful framework for aligning LLMs with human-centric evaluation criteria in the context of ATS. By constructing reward models through annotated datasets that encode preferences for conciseness, factual accuracy, and coherence, RLHF enables iterative policy optimization to refine summary generation. This approach addresses critical challenges in abstractive summarization, such as information compression while preserving key semantic elements, elimination of redundant expressions, and maintenance of factual consistency across multi-sentence outputs. [252] systematically replicated and extended the RLHF-based scaling behaviors reported in OpenAI’s TL;DR summarization work by constructing a novel RLHF pipeline from scratch, meticulously enumerating over 20 critical implementation details, and documenting key insights throughout the reproduction process. [253] proposed a novel methodology to reduce the dependency on extensive human preference annotations in RLHF by infusing domain-specific knowledge into the reward model, thereby enabling task-specific alignment with minimal annotated data while maintaining interpretability and adaptability to diverse human value preferences.
7.3.2 External adapters fine-tuning
@@
Training external adapters allows models to learn the distribution of target data without modifying the original parameters of the LLM. [254] proposed an efficient few-shot method using adapters, where the adapters were pre-trained on a large corpus and fine-tuned on a smaller human-annotated dataset. Similarly, [255] introduced a contrastive learning approach for supervised ATS, aiming to maximize the similarity between a document, its gold-standard summary, and model-generated summaries. Additionally, [256] combined encoder and decoder foundation models into a single model, AdaMo, and implemented adaptive knowledge transfer techniques such as continuous pretraining and intermediate fine-tuning, along with task-specific designs for sequence-to-sequence learning.
LoRA (Low-Rank Adaptation)[257] has been applied in ATS to efficiently fine-tune LLMs by introducing low-rank matrices that adapt pre-trained weights without modifying the original architecture, significantly reducing computational and memory costs. This method enables models to maintain high summarization performance while being optimized for specific tasks or domains, such as biomedical or technical text, with minimal additional parameters. The DataHacks team developed a perspective-aware summarization approach[258] for healthcare community question-answering forums by fine-tuning LLMs with various LoRA configurations to balance performance and computational efficiency. Their method involved identifying and classifying perspective-specific spans, followed by generating summaries based on distinct perspectives using adapted prompt strategies. [259] investigated the efficacy of fine-tuning open-source LLMs for aspect-based summarization by adapting pre-trained models such as Llama2, Mistral, Gemma, and Aya to domain-specific tasks. They employed supervised fine-tuning using prompt-completion pairs from the OASUM dataset, combined with parameter-efficient techniques like Quantized Low-Rank Adaptation (QLoRA) and PEFT to enhance training efficiency and model specialization for aspect-focused information extraction.
Fine-tuning LLMs for ATS offers significant advantages, including enhanced domain-specific adaptation to specialized contexts (e.g., biomedical or legal texts) by capturing nuanced linguistic patterns and task requirements. Techniques like LoRA enable parameter-efficient training, reducing computational costs while maintaining performance, and mitigating the need for extensive manual annotation through knowledge transfer from pre-training. However, this approach faces challenges such as overfitting to limited training data in low-resource domains, high computational demands for training and inference, and potential inheritance of biases or factual inaccuracies from training datasets. Domain mismatches between pre-training objectives and fine-tuning tasks may further degrade performance when source texts differ structurally or semantically from the target domain. Additionally, the “black-box” nature of LLMs complicates interpretability, posing barriers for high-stakes applications where transparency is critical. These trade-offs highlight the need for balanced strategies to optimize domain-specific performance while addressing computational, ethical, and interpretability constraints in ATS deployment.
7.4 Knowledge Distillation from LLMs for ATS
@@
A knowledge distillation system consists of three key components: knowledge, a distillation algorithm, and a teacher-student architecture [297]. In this process, a smaller student model is supervised by a larger teacher model [298]. For ATS, this typically involves an offline distillation process where an LLM serves as a fixed teacher model, with its outputs providing additional supervision to train a smaller student model. This approach enables the student model to inherit the LLM’s robust summarization capabilities, making it particularly useful in scenarios with limited computational resources or strict data privacy requirements. As shown in Table 10, [220] improved the quality of aspect-triple rationales and summaries through a dual scoring system, followed by a curriculum-learning strategy to train a smaller local model on ATS tasks. Similarly, [260] addressed low-resource cross-lingual summarization using mBART for incremental training and reinforcement learning to optimize discrete prompts. [235] proposed a three-step method for adapting smaller models to summarize forum discussions by sampling from a large corpus, retrieving annotated prompts, and filtering low-quality data.
@@
Knowledge distillation from LLMs for ATS aims to transfer summarization capabilities from complex teacher models to smaller student models, enabling efficient deployment while retaining key strengths like factual consistency and coherence. This approach reduces computational costs and allows real-time applications on resource-limited devices, leveraging soft label distillation to mimic the teacher’s nuanced outputs. However, distilled models inherit biases or errors from the teacher, struggle with multi-step reasoning tasks, and often underperform in summarization compared to their larger counterparts. While cost-effective and scalable, their effectiveness depends on careful mitigation of inherited flaws and domain-specific fine-tuning, requiring a balance between efficiency and accuracy.
@@ 8.1 Metrics based on Term Overlap
ROUGE (Recall-Oriented Understudy for Gisting Evaluation) [301] is a widely used metric for ATS evaluation, measuring overlap between candidate summaries ($C$) and reference summaries ($S$). ROUGE-N calculates n-gram overlap as:
@@
METEOR (Metric for Evaluation of Translation with Explicit ORdering) [303] was developed to overcome the rigidity of BLEU, which relies on exact n-gram matches. METEOR evaluates summaries by aligning words between a candidate and reference summary using flexible matching techniques, including synonyms, stemming, and paraphrases, facilitated by resources like WordNet. This allows METEOR to account for linguistic variations and semantic similarities that BLEU cannot capture. Additionally, METEOR differentiates between function words (e.g., “the”, “and”) and content words (e.g., “run”, “house”), giving more weight to content words to better reflect meaning. By incorporating these features, METEOR provides a more nuanced and flexible evaluation metric for summarization tasks.
@@ 8.2 Similarity-based ATS Metrics
Similarity-based metrics assess summary quality by measuring the semantic similarity between the candidate summaries and either the reference summaries or the source document. These metrics can be broadly grouped into two types: reference-based metrics, which compare the candidate summary against ground-truth references (e.g., BERTScore, BARTScore, BLEURT), and source-based metrics, which assess consistency or entailment between the candidate and the source document (e.g., FactCC, SummaC).
BERTScore [305] evaluates text quality by computing a similarity score between each token in the candidate summary and each token in the reference summary, using contextual embeddings from pre-trained language models like BERT. These approaches capture semantic meaning and contextual relationships, offering a more flexible and effective evaluation compared to traditional n-gram-based metrics.
BARTScore [306] evaluates text quality by computing the generation likelihood of a candidate summary conditioned on a reference summary or source document, using a pre-trained sequence-to-sequence language model like BART. By treating evaluation as a text generation task, BARTScore leverages the full generative capacity of large models to capture semantic adequacy and fluency. This approach enables more nuanced assessment of summary quality, especially in abstractive scenarios, and has been shown to correlate well with human judgments, outperforming traditional n-gram and embedding-based metrics.
BLEURT [307] builds on BERT embeddings by incorporating additional fine-tuning on human-annotated data and synthetic examples. This fine-tuning process allows BLEURT to explicitly model text quality judgments, aligning more effectively with task-specific requirements such as fluency, coherence, and relevance. Compared to BERTScore, BLEURT provides a more comprehensive evaluation by integrating various aspects of text quality.
@@
$\text{BERT-}i\text{BLEU}$ $=\left(\frac{\beta*\text{BERT-}score^{-1}+1.0*(1-self\text{-BLEU})^{-1}}{\beta+1.0}\right)^{-1}$ (9)
$self\text{-BLEU}$ $=\text{BLEU}(source,candidate)$
@@
FactCC [169] takes a source document and a summary statement as input and outputs the fact consistency label for that summary statement. FactCC is a weakly-supervised model proposed to verify factual consistency in summaries by predicting consistency, extracting supporting spans, and identifying inconsistent spans, which outperforms strongly supervised models and aids human verification. By modeling different types of fact consistency errors, the method achieves improved performance consistency evaluation.
@@ 8.3 LLM-based ATS Metrics
Traditional evaluation metrics for assessing summary quality primarily focus on specific dimensions such as lexical overlap (e.g., ROUGE) or semantic similarity (e.g., BERTScore). However, these approaches exhibit notable limitations in capturing the multifaceted nature of summarization tasks. By contrast, large language models (LLMs) demonstrate a capacity for more nuanced and comprehensive analyses through prompt engineering that leverages their contextual understanding and reasoning capabilities. Inspired by [314], existing LLM-based ATS metrics can be systematically categorized into four following groups:
8.3.1 Score-based LLM Evaluation Metrics
Score-based evaluation metric prompts LLMs to generate continuous quality scores for summaries, typically on a predefined scale. For example, LLMs might evaluate fluency, coherence, or relevance directly by assigning numerical ratings based on the input text and reference (if provided).
G-EVAL [315] introduced a human-aligned reference-free framework that leverages GPT-4 with chain-of-thought reasoning and form-filling paradigms to assess generated summaries, addressing limitations of traditional reference-based metrics in capturing creativity and diversity.
ICE [316] leverages few‑shot in‑context learning. By exposing GPT‑3 to four sets of manually scored examples, the model is prompted to output dimension‑specific scores (ranging from 0 to 1) for target summaries, enabling the assessment of coherence, relevance, fluency, and consistency without additional training.
SumAutoEval [317] decomposed summarization quality into four granular, entity-centric dimensions: Completeness, Correctness, Alignment, and Readability, by extracting fine-grained entities from both the ground-truth and candidate summaries and then prompting an LLM with a suite of consistency-checked prompts to verify each entity’s coverage, factual accuracy, positional relevance, and linguistic fluency, finally aggregating the entity-level judgments into interpretable, human-aligned scalar scores.
OP-I-PROMPT [318] distilled the evaluation of opinion summaries into a single, dimension-independent prompt that instructed an LLM to score each summary on seven discrete aspects: fluency, coherence, relevance, faithfulness, aspect coverage, sentiment consistency, and specificity.
CREAM [319] distills completeness, conciseness, and faithfulness into a single reference‑free, comparison‑based prompt for GPT‑4, instructing it to extract key facts from paired meeting summaries and score each summary on these three aspects without relying on gold references or source transcripts.
Talec [320] were operationalized through a set of ten binary labels: five unacceptable labels (score = 0) and five acceptable labels (score = 1), that jointly determined a final score of 0, 1, or 2 (full). These labels assessed whether a response failed to meet requirements, contained incorrect or unrelated content, refused to answer, exhibited untranslated text, confusing structure, external links, stiffness, repetition, subject imprecision, or incompleteness.
Score-based evaluation metrics leverage the reasoning capabilities of large language models (LLMs) to generate continuous quality scores for summaries, typically on a predefined scale. This approach enables direct assessment of dimensions such as fluency, coherence, and relevance by assigning numerical ratings based on input text and reference summaries (if available). The primary strength of this method lies in its quantitative precision, allowing for straightforward comparisons across models or outputs. However, the reliance on predefined criteria introduces subjectivity, as the interpretation of “high fluency” or “strong relevance” may vary across LLMs or evaluators. Additionally, this method often depends heavily on high-quality reference summaries, which are not always accessible or representative of diverse user preferences, potentially limiting its applicability in real-world scenarios.
8.3.2 Probability-based LLM Evaluate Metrics
Probability-based evaluation metrics compute the likelihood of generating a summary given a source text or reference. This method assumes that higher-generation probabilities correlate with better quality, reflecting semantic alignment and fluency.
GPTScore [321] prompted LLM evaluator to calculate the generation probability of generated text with definitions of evaluation tasks and aspects that leveraged the zero-shot reasoning capabilities of LLMs to assess generated summaries through natural language instructions.
ProbDiff [322] proposed a self‑evaluation metric that quantifies an LLM’s competence by computing, for each query, the negative log‑probability difference between its initial response and a K‑step refined response. Under the empirically supported assumption that stronger models exhibit flatter probability distributions, a smaller discrepancy indicates higher confidence and better performance, enabling model evaluation without external annotations or additional training.
BWRS [323] proposed a Bayesian calibration framework that adjusts win‑rate estimates from imperfect LLM evaluators by modeling both observed outcomes and evaluator accuracy as Beta‑Bernoulli processes. By drawing samples from the posterior distributions of evaluator true positive or negative rates and observed win counts, BWRS derives an unbiased estimate of the true model win rate. This calibration both corrects evaluator bias and quantifies uncertainty, enabling more reliable pairwise comparisons without extensive human labeling.
Probability-based evaluation metrics assess summary quality by computing the likelihood of generating a summary given a source text or reference, under the assumption that higher-generation probabilities correlate with better alignment and fluency. This method capitalizes on the pretraining of LLMs on vast corpora, enabling objective probabilistic assessments grounded in statistical patterns. A key advantage is its efficiency, as it requires minimal computational overhead compared to human annotation. However, this approach risks conflating fluency with factual accuracy, as a grammatically correct but semantically incorrect summary may receive a high probability score. Furthermore, the method may prioritize common patterns in training data over novel or user-specific summarization needs, introducing biases that compromise its reliability in diverse contexts.
8.3.3 LLM-based Pairwise Comparison Metrics
Pairwise comparison metrics compare two summaries to determine which one is superior in terms of quality, coherence, or relevance. This method is particularly useful for ranking models or identifying strengths/weaknesses in competing outputs.
BOOOOKSCORE [324] constructed a reference-free coherence metric by distilling human-annotated coherence-error taxonomies into a GPT-4 driven, few-shot prompt that classified each summary sentence against eight error types and reported the proportion of unconfused sentences, thus enabling scalable, LLM-agnostic evaluation of book-length summarization systems.
Calibrated Confidence Score [325] computes the geometric mean of per-token generation probabilities and then applies Platt scaling using human‑similarity labels (derived from BERTScore) to produce a calibrated confidence score that aligns model confidence with human judgments.
PRE [326] introduced the Peer Review Evaluator, which quantifies alignment between automated and human judgments by computing agreement rate, Kendall’s tau, and Spearman correlation across summarization and non-factoid question‑answering tasks.
FineSurE [327] decomposed summarization evaluation into three fine-grained dimensions: faithfulness, completeness, and conciseness, by prompting a large language model (1) to classify each summary sentence into nine factuality-error types and (2) to align a set of human- or LLM-extracted keyfacts with the corresponding summary sentences, then aggregating sentence- and keyfact-level judgments into interpretable percentage scores that correlated strongly with human annotations.
ACUEval [328] decomposed each summary into atomic content units and validated them against the source document via an LLM-driven two-step pipeline, yielding interpretable, fine-grained faithfulness scores that improved balanced accuracy over prior metrics by 3% and enabled actionable error correction.
Pairwise comparison metrics evaluate two summaries to determine relative superiority in terms of quality, coherence, or relevance, mimicking human judgments by ranking outputs based on their strengths. This method is particularly effective for model benchmarking and identifying strengths/weaknesses in competing outputs, as it aligns with how humans intuitively compare alternatives. Its flexibility allows application even without reference summaries, relying solely on the LLM’s contextual reasoning. Nevertheless, this approach lacks an absolute benchmark, making it challenging to quantify overall performance improvements. Additionally, the absence of explicit criteria for comparison (e.g., prioritizing factual accuracy over conciseness) can lead to inconsistent evaluations, especially when summaries differ significantly in multiple dimensions.
8.3.4 Hybrid Evaluation Metrics
Hybrid evaluation metrics use multiple metrics or LLMs to assess summaries from different perspectives. This reduces bias and captures a wider range of quality dimensions.
Humanlike [329] examined four human-inspired evaluation protocols, including Likert-scale rating, pairwise comparison, Pyramid content-unit matching, and binary factuality verification, through which ChatGPT assessed summarization quality on five benchmark corpora. By computing Spearman correlations and accuracy against expert annotations, the authors demonstrated that ChatGPT’s judgments, particularly when prompted with concise task descriptions, surpassed traditional surface-level metrics (ROUGE, BERTScore, BARTScore) and rivaled human inter-annotator agreement, yielding a cost-effective, reproducible alternative for summary evaluation.
DRPE [330] leveraged a diverse set of LLM-based role-players, statically defined for objective dimensions such as coherence and grammar and dynamically generated for subjective dimensions such as interestingness,to conduct pairwise comparisons between machine-generated and reference summaries, aggregated the resulting votes via batch prompting, and was empirically shown to correlate more strongly with human judgments than conventional overlap- or embedding-based measures.
FENICE [331] introduced an interpretable and computationally efficient factuality metric for summarization evaluation that extracted atomic claims from summaries via an LLM and distilled them into a T5-base generator, aligned each claim with document spans at sentence, paragraph and document granularity through a DeBERTa-v3-large NLI model, refined the entailment scores with coreference resolution, and achieved state-of-the-art balanced accuracy on the AGGREFACT benchmark while surpassing baselines on a newly annotated long-form summarization dataset.
DEBATE [332] (Devil’s Advocate-Based Assessment and Text Evaluation), a multi-agent LLM framework that instantiated a Commander–Scorer–Critic architecture in which a Critic was explicitly prompted to adopt the role of Devil’s Advocate and iteratively challenged the Scorer’s judgements. By enabling structured debates capped at a pre-defined number of rounds, DEBATE alleviated single-agent biases, yielded scores that exhibited significantly higher correlations with human ratings on the SummEval and Topical-Chat meta-evaluation benchmarks, and established a new state-of-the-art among reference-free, LLM-based NLG evaluators.
Hybrid evaluation metrics integrate multiple methods, such as score-based, probability-based, and pairwise comparisons, or combine outputs from different LLMs to assess summaries from diverse perspectives. This approach mitigates the limitations of individual metrics by aggregating their strengths, such as reducing subjectivity in scoring or capturing both absolute and relative quality dimensions. Its adaptability allows customization for specific tasks, such as emphasizing factual accuracy in news summarization or conciseness in technical documents. However, hybrid methods introduce complexity in design and interpretation, as aggregating diverse outputs (e.g., scores, probabilities, rankings) may obscure insights into specific quality aspects. Moreover, the resource intensity of integrating multiple models or metrics increases computational costs, necessitating careful calibration to balance comprehensiveness and efficiency.